Question 2.6 [3, 3, 3]



Suppose that the height of seven-year old children is a normal random variable with mean 1.1m and



standard deviation 0.1m.



a) What proportion of seven-year old children are shorter than 0.95m?



b) What is the probability that a seven-year old child is taller than 1m but shorter than 1.175m?



c) Suppose that a family restaurant has a children’s play-place. The facilities in the play-place



can only be used by children below a certain height. What should the cut-off height be to



ensure that 99% of seven-year olds have access to the play-place?


1
Expert's answer
2022-06-03T03:16:34-0400

a)


"P(X<0.95)=P(Z<\dfrac{0.95-1.1}{0.1})"

"=P(Z<-1.5)=0.0668"

b)


"P(1<X<1.175)=P(Z<\dfrac{1.175-1.1}{0.1})"

"-P(Z\le\dfrac{1-1.1}{0.1})=P(Z<0.75)-P(Z\le-1)"

"=0.6147"

c)


"P(X<x)=P(Z<\dfrac{x-1.1}{0.1})=0.99"

"\dfrac{x-1.1}{0.1}=2.3263"




"x=0.23263+1.1"

"x=1.333\ m"